OpenStudy (anonymous):

the volume of a cylinder is 328cm3. find the radius of the base if the cylinder has a height of 8cm. round your answer to the nearest 0.1cm

OpenStudy (anonymous):

Volume is base times height. You have volume and height, so volume divided by height is base. Once you have the area of the base, since A=pi*r^2 solve for r: \[r=\sqrt{\frac{A}{\pi}}\] Substituting: A=328/8 then divide that by pi finally take the square root and you're done.

OpenStudy (anonymous):

can u explain a little more?

OpenStudy (anonymous):

Sure. Normally you get the base info and you calculate the area, then you multiply by height and you get the volume. In this case They give you volume and they expect you to get the radius, so you work backwards. EAch formula has to be rearranged. So you have to start at the end. The end is V=base area times height first you solve for base area, because you have the other two. If V=A*h then you move the h dividing to the side with the V you get A=V/h That's where I did A=328/8 Next the area of a circle is: \[A=pi*r^2\] to solve for r first move the pi. Since it's multiplying the r here, then you do the opposite on the other side and move it dividing. \[r^2=\frac{A}{pi}\] that would give you r squared, so you do the opposite on the other side (square root) and you're done. \[r=\sqrt{\frac{A}{\pi}}\]